OpenStudy (anonymous):

rationalize the square root √64 over √2

OpenStudy (lncognlto):

What do you think the answer is?

OpenStudy (anonymous):

4√2

OpenStudy (lncognlto):

Correct

OpenStudy (lncognlto):

=8/√2 =8√2/2 =4√2